find the quotient and remainder of f(x) divided by d(x): f(x) = 3x^5 - 4x^3 + x + 5 d(x) = x^2 - 2x + 3
Answer by sachi(548)

f(x) = 3x^5 - 4x^3 + x + 5
d(x) = x^2 - 2x + 3
f(x) divided by d(x)=
x^2 - 2x + 3 )3x^5 - 4x^3 + x + 5 (3x^3+6x^2-x-20
3x^5-6x^4+9x^3
- + -
------------------
6x^4-13x^3+x+5
6x^4-12x^3+18x^2
- + -
------------------
-x^3-18x^2+x+5
-x^3+2x^2-3x
+ - +
------------
-20x^2+4x+5
-20x^2+40x-60
+ - +
-------------
-36x+65
----------
so the quotient = 3x^3+6x^2-x-20 & the reminder=-36x+65
ans